# A Survey of LLM-Driven AI Agent Communication: Protocols, Security Risks, and Defense Countermeasures

### v3

@@ VI-C Defense Countermeasure Prospect
We will discuss the possible defense countermeasures targeting the proposed security risks caused by compromised agents (against not only benign agents but also benign users and environments). As a result, we hope our work can motivate more discussion on this area and benefit the future design/deployment of agent communication.
VI-C1 Countermeasures for CS-based Communication Risks
To mitigate the risks summarized in Section VI-B1, we hope developers to achieve the following strategies/mechanisms.
@@ VI-C1 Countermeasures for CS-based Communication Risks
Registration Verification and Monitoring. To mitigate registration pollution, agent servers need to build a strict registration access mechanism using techniques like zero-trust authentication [257] to verify the registration of an agent. Besides, servers should monitor the dynamic behaviors at the agent-level and IP-level. For example, the number of registrations for each IP address should be limited, and frequent registration/deregistration should be treated as abnormal behavior. Once malicious registration is detected, automatic interception is immediately triggered, and suspicious agents/IPs are added to the blacklist. Syros et al. [260] propose SAGA. SAGA makes users register agents with the central entity Provider and implement fine-grained interaction control using encrypted access control tokens, thereby balancing security and performance.
@@
Anti-manipulation Optimization. To mitigate SEO poisoning, agent servers should deploy robust agent searching algorithms. For example, they can introduce adversarial training to enhance the model’s anti-manipulation ability, or conduct semantic blurring/replacing on search keywords, to prevent malicious agents from improving rankings. Besides, the search algorithms can deploy a random factor to ensure a ratio of randomly selected agents in the final list. Meanwhile, dynamically updating parameters and inducing historical response quality are also helpful.
VI-C2 Countermeasures for P2P-based Communication Risks
@@
Task Lifecycle Monitoring. We think the non-convergence problem is stubborn and hard to eliminate as long as the P2P architecture is not changed fundamentally. As a result, the method of mitigating this problem is to monitor the task lifecycle. Each access point should deploy a coordinator. For agent-agent communication, this coordinator monitors the execution status. When it detects that the task interaction is trapped in a loop (e.g., no progress after N consecutive rounds of responses) or the communication time exceeds a threshold, it forcibly terminates the non-convergent communication. At the same time, the abnormal patterns and the communication participants are recorded for further analysis. He et al. [105] propose a Trust Management System (TMS), which deploys message-level and agent-level trust evaluation. TMS can dynamically monitor agent communication, execute threshold-driven filtering strategies, and achieve agent-level violation record tracking. Zhang et al. [336] propose G-Memory, a hierarchical memory system. G-Memory manages the interaction history of agent communication through three-layer graph structures of Insight Graph, Query Graph, and Interaction Graph, thereby achieving the evolution of the agent team. Ebrahimi et al. [66] propose an anti-adversarial multi-agent system based on Credibility Score. It models query answering as an iterative cooperative game, distributes rewards through Contribution Score, and dynamically updates the credibility of each agent based on historical performance.
End-to-end Encryption Enhancement. Although some existing protocols like A2A and ANP support end-to-end encryption and integrity verification mechanisms, the risks of MITM attacks are not eliminated due to various deployment errors or protocol vulnerabilities. As a result, besides deploying such security algorithms, the community should also adopt other strategies to enhance the end-to-end communication, such as timely updating versions to fix bugs and designing a transmission path redundancy mechanism. For example, Sharma et al. [245] point out that using encrypted communication is necessary to enhance the security of A2A. We believe it is a long-term process to defend against MITM attacks.
VI-C3 Countermeasures for Universal Risks
Identity Authentication. The identity authentication of agents is critical to defending against agent spoofing in multi-agent systems. Sharma et al. [245] also emphasize the importance of authentication in deploying the A2A protocol. As we have analyzed, identity authentication may show better performance in the CS-based communication if capability verification is deployed at the same time. In contrast, for P2P-based communication, authentication can mitigate agent spoofing caused by MITM attacks, but will fail if the attackers have a legal identity but exaggerated capability descriptions. Since P2P-based communication inherently lacks the ability to verify the capability of agents, we think agent spoofing may still exist for a long time. Shah et al. [243] ensure the immutability of online transactions through blockchain, use multi-factor authentication (MFA) for identity verification, and rely on a machine-learning-based anomaly detection system to identify abnormal transactions in real-time.
Agent Behavior Auditing and Accountability. To avoid agent exploitation/Trojan, agent bullying, and responsibility evasion, it is necessary to audit the behaviors of agents to avoid damage/influences to the task execution. For example, there should be a logging mechanism that periodically records the communication contents, and AI algorithms to dynamically calculate the responsibility of each action. Rastogi et al. propose AdaTest++, allowing humans and AI to collectively audit the behaviors of LLMs [232]. Amirizaniani et al. [13] propose a multi-probe method to detect potential issues such as bias and hallucinations caused by LLMs. Mokander et al. [205] design a three-layered approach, auditing LLMs using governance audits, model audits, and application audits. Das et al. [52] propose CMPL, which generates probes through LLM and combines with human verification, adopts sub-goal-driven and reactive strategies, and audits the privacy leakage risks of agents from both explicit and implicit aspects. Jones [131] proposes a series of systems to detect rare failures, unknown multimodal system failures, and LLM semantic biases, respectively. Nasim et al. [208] propose a Governance Judge Framework. By deploying input aggregation, evaluation logic, and a decision-making module, it realizes the automated monitoring of agent communication to address issues such as performance monitoring, fault detection, and compliance auditing. Deshpande et al. [58] propose the TRAIL dataset containing 148 manually annotated traces, and use it to evaluate the LLM’s ability to analyze agent workflow traces. Although existing studies can provide valuable insights, the research on agent behavior auditing still needs long-term efforts. Tamang et al. [261] propose the Enforcement Agent (EA) framework, which embeds supervisory agents in a multi-agent system to achieve real-time monitoring, detection of abnormal behaviors, and intervention of other agents. Toh et al. [271] proposes the Modular Speaker Architecture (MSA). By decomposing dialogue management into three core modules: Speaker Role Assignment, Responsibility Tracking, and Contextual Integrity, and combining with the Minimal Speaker Logic (MSL) to formalize responsibility transfer, MSA addresses the issues of accountability in multi-agent systems. Fan et al. [71] propose PeerGuard, which uses a mutual reasoning mechanism among agents to detect the inconsistencies in other agents’ reasoning processes and answers, thereby identifying compromised agents. Jiang et al. [129] propose Thought-Aligner, which uses a model trained with contrastive learning to real-time correct high-risk thoughts before the agent executes actions, thereby avoiding the dangerous behaviors of agents.
Access Control. To mitigate privacy leakage, the access control among agents is a core component for the future agent ecosystem. Although end-to-end encryption can avoid sniffing from external attackers to some extent, it cannot mitigate the unintentional privacy leakage among agents. Access control should assign access permission tags to different agents and ensure that agents need to attach permission proofs when communicating. In this way, agents with low-level permissions cannot obtain the high-level sensitive information from other agents. Zhang et al. [340] design the AgentSandbox framework, which uses the separation of persistent agents and temporary agents, data minimization, and I/O firewalls, realizing the security of agents in solving complex tasks. Kim et al. [140] propose the PFI framework, which defends against authority-related attacks through three major technologies: agent isolation, secure untrusted data processing, and privilege escalation guards. Wang et al. [279] propose AgentSpec. It allows users to define rules containing trigger events, predicate checks, and execution mechanisms through a domain-specific language to ensure the safety of agent behavior.
@@ VI-C3 Countermeasures for Universal Risks
Multi-Source Channel Isolation. In multi-agent settings, input isolation is critical to prevent malicious intent from propagating between agents. Systems should avoid concatenating raw messages from other agents and instead extract structured key information while stripping control-oriented content. Furthermore, deploying a safety coordination agent to review, sanitize, or flag inter-agent messages can effectively mitigate the potential attack propagation within multi-agent systems.
Attack Modeling and Testing. To discover unknown vulnerabilities, designing an attack generation testing framework is also an effective method. By generating different attack vectors to the target agent system, developers can find new loopholes based on abnormal outputs. Gandhi et al. [78] propose the ATAG framework. By extending the MulVAL tool [215], introducing custom facts and interaction rules, and combining with the newly constructed LLM Vulnerability Database (LVD), ATAG realizes the modeling and analysis of attacks against multi-agent scenarios, such as privacy leakage. Yu et al. [326] propose NetSafe, which models the multi-agent network as a directed graph. NetSafe combines three types of attack strategies: error information injection, bias induction, and harmful information eliciting. It evaluates topological security through static and dynamic metrics.
Agent Orchestration. To avoid task flooding or DoS attacks against the agent-agent communication, achieving agent orchestration is necessary. It can automatically optimize the task scheduling and assigning process to reduce the communication overhead, and can also optimize the prompts generated by agents to save computing resources for the involved agents. How et al. [112] propose HALO. HALO realizes dynamic task decomposition and role generation through a three-layer collaborative architecture. It uses Monte Carlo Tree Search to explore the optimal reasoning trajectory and transforms user queries into task-specific prompts through the adaptive prompt refinement module. Owotogbe [216] designs a chaos engineering framework in three stages (conceptual framework, framework development, empirical verification). By simulating interference scenarios such as agent failures and communication delays, and combining multi-perspective literature reviews and GitHub analysis, this work aims to systematically identify vulnerabilities and enhance the resilience of agent systems.
@@ VI-D Takeaways
In this section, we categorize two major agent-agent communication architectures: CS-based and P2P-based. Accordingly, the security risks are also multifaceted: CS-based architecture puts heavy burdens on the centralized agent servers, such as registration pollution and SEO poisoning. P2P-based architecture suffers from the lack of efficient and centralized management of agents, such as non-convergence and man-in-the-middle attacks. Besides, both of them are vulnerable to universal risks, such as agent spoofing, bullying, and privacy leakage. We also discussed potential defense countermeasures targeting each risk. The defenses discussed in this section can also mitigate the risks from compromised agents against user (Section V-C) and environments (Section VII-C). We believe that as agent-agent communication continues to grow, more vulnerabilities in this process will be discovered.

### v4

@@ VI-C Defense Countermeasure Prospect
We will discuss the possible defense countermeasures targeting the proposed security risks associated with three categories of communication protocols (universal, client-server, and peer-to-peer). As a result, we hope our work can motivate more discussion on this area and benefit the future design/deployment of agent communication.
VI-C1 Defenses on Data Transmission Layer (L1)
The defenses are classified based on the risks from L1.
(i) Against Universal Risks in all Communication Modes. As we analyzed in Section VI-B1, the reason for risks on this layer is the same as in user-agent communication. Therefore, the defense countermeasures are the same as Section V-C1.
(ii) Against Unique Risks in CS-based Communication.
D1: Controller Isolation and Enhancement. To secure the centralized controller in CS-based communication, multi-agent systems must enforce strict protection over the controller’s data transmission. First, all inter-agent messages routed through the controller should be authenticated, integrity-checked, and rate-limited. Second, the controller must operate in an isolated and hardened execution environment (e.g., sandboxed runtime, privilege separation) to minimize the impact of compromise. Third, maintaining a redundant or failover controller, combined with state replication or checkpoint synchronization, can prevent single-point-of-failure outages. Finally, continuous anomaly detection—such as monitoring abnormal routing patterns, message-drop rates, or suspicious rewrite behaviors—can provide early warning of controller hijacking attempts. These measures collectively prevent the centralized controller from becoming an attack amplifier for the entire system.
VI-C2 Defenses on Interaction Protocol Layer (L2)
The defenses are classified based on the risks from L2.
(i) Against Universal Risks in all Communication Modes.
D1: Identity Authentication and Capability Verification. The identity authentication of agents is critical to defending against agent spoofing in multi-agent systems. Sharma et al. [sharma2024secure] also emphasize the importance of authentication in deploying the A2A protocol. As we have analyzed, identity authentication may show better performance in the CS-based communication if capability verification is deployed at the same time. In contrast, for P2P-based communication, authentication can mitigate agent spoofing caused by MITM attacks, but will fail if the attackers have a legal identity but exaggerated capability descriptions. Since P2P-based communication inherently lacks the ability to verify the capability of agents, we think agent spoofing may still exist for a long time. Shah et al. [shah2025enforcing] ensure the immutability of online transactions through blockchain, use multi-factor authentication (MFA) for identity verification, and rely on a machine-learning-based anomaly detection system to identify abnormal transactions in real-time. Beyond transaction-level checks, LLM fingerprinting serves as a critical authentication layer to verify model integrity and identity, spanning comprehensive surveys and evaluations [xu2025copyright, naseryAreRobustLLM2025], embedding mechanisms [wangFPEditRobustLLM2025, naseryScalableFingerprintingLarge2025, gloaguenLLMFingerprintingSemantically2025, xuEverTracerHuntingStolen2025, xuCTCCRobustStealthy2025, yuePREEHarmlessAdaptive2025, xu2025insty, wuEditMFDrawingInvisible2025, wuImFImplicitFingerprint2025], erasure defenses [zhangMEraserEffectiveFingerprint2025], transferability analyses [xuUnlockingEffectivenessLoRAFP2025, xu2025fingerprintvectorenablingscalable], and generation methods [xu2025rapsmrobustadversarialprompt, pasquiniLLMmapFingerprintingLarge2025, shaoReadingLinesReliable2025, wuLLMDNATracing2025, zhangREEFRepresentationEncoding2025].
D2: Behavior Auditing and Responsibility Tracing. To avoid agent exploitation/Trojan, agent bullying, and responsibility evasion, it is necessary to audit the behaviors of agents to avoid damage/influences to the task execution. For example, there should be a logging mechanism that periodically records the communication contents, and AI algorithms to dynamically calculate the responsibility of each action. Rastogi et al. propose AdaTest++, allowing humans and AI to collectively audit the behaviors of LLMs [rastogi2023supporting]. Amirizaniani et al. [amirizaniani2024auditllm] propose a multi-probe method to detect potential issues such as bias and hallucinations caused by LLMs. Mokander et al. [mokander2024auditing] design a three-layered approach, auditing LLMs using governance audits, model audits, and application audits. Das et al. [das2025disclosure] propose CMPL, which generates probes through LLM and combines with human verification, adopts sub-goal-driven and reactive strategies, and audits the privacy leakage risks of agents from both explicit and implicit aspects. Jones [jones2025scalable] proposes a series of systems to detect rare failures, unknown multimodal system failures, and LLM semantic biases, respectively. Nasim et al. [nasim2025governance] propose a Governance Judge Framework. By deploying input aggregation, evaluation logic, and a decision-making module, it realizes the automated monitoring of agent communication to address issues such as performance monitoring, fault detection, and compliance auditing. Deshpande et al. [deshpande2025trail] propose the TRAIL dataset containing 148 manually annotated traces, and use it to evaluate the LLM’s ability to analyze agent workflow traces. Although existing studies can provide valuable insights, the research on agent behavior auditing still needs long-term efforts. Tamang et al. [tamang2025enforcement] propose the Enforcement Agent (EA) framework, which embeds supervisory agents in a multi-agent system to achieve real-time monitoring, detection of abnormal behaviors, and intervention of other agents. Toh et al. [toh2025modular] proposes the Modular Speaker Architecture (MSA). By decomposing dialogue management into three core modules: Speaker Role Assignment, Responsibility Tracking, and Contextual Integrity, and combining with the Minimal Speaker Logic (MSL) to formalize responsibility transfer, MSA addresses the issues of accountability in multi-agent systems. Fan et al. [fan2025peerguard] propose PeerGuard, which uses a mutual reasoning mechanism among agents to detect the inconsistencies in other agents’ reasoning processes and answers, thereby identifying compromised agents. Jiang et al. [jiang2025think] propose Thought-Aligner, which uses a model trained with contrastive learning to real-time correct high-risk thoughts before the agent executes actions, thereby avoiding the dangerous behaviors of agents.
D3: Causal Tracing and Localization. Beyond discovering unknown vulnerabilities, attack-generation testing can also help mitigate responsibility-evasion issues by exposing which agents and interaction steps contribute to harmful outcomes. Frameworks such as ATAG [gandhi2025atag] and NetSafe [yu2024netsafe] generate diverse adversarial perturbations and analyze their propagation through multi-agent workflows. By modeling the system as a causal interaction graph and evaluating how injected errors or biases spread, these tools enable developers to identify the specific agent or message that triggers downstream deviations. Integrating lightweight provenance logging or causal-path tracing into such testing frameworks further strengthens accountability, making it easier to attribute failures to misbehavior, errors, or malicious modifications by individual agents.
D4: Defense for Denial of Service. To avoid DoS attacks against the agent-agent communication, achieving agent orchestration is necessary. It can automatically optimize the task scheduling and assigning process to reduce the communication overhead, and can also optimize the prompts generated by agents to save computing resources for the involved agents. How et al. [hou2025halo] propose HALO. HALO realizes dynamic task decomposition and role generation through a three-layer collaborative architecture. It uses Monte Carlo Tree Search to explore the optimal reasoning trajectory and transforms user queries into task-specific prompts through the adaptive prompt refinement module. Owotogbe [owotogbe2025assessing] designs a chaos engineering framework in three stages (conceptual framework, framework development, empirical verification). By simulating interference scenarios such as agent failures and communication delays, and combining multi-perspective literature reviews and GitHub analysis, this work aims to systematically identify vulnerabilities and enhance the resilience of agent systems.
(ii) Against Unique Risks in CS-based Communication.
D5: Defense for Registration Pollution. To address the registration pollution risks that arise in CS-based communication, we propose the following defense strategies.
@@ VI-C2 Defenses on Interaction Protocol Layer (L2)
Registration Verification and Monitoring. To mitigate registration pollution, agent servers need to build a strict registration access mechanism using techniques like zero-trust authentication [stafford2020zero] to verify the registration of an agent. Besides, servers should monitor the dynamic behaviors at the agent-level and IP-level. For example, the number of registrations for each IP address should be limited, and frequent registration/deregistration should be treated as abnormal behavior. Once malicious registration is detected, automatic interception is immediately triggered, and suspicious agents/IPs are added to the blacklist. Syros et al. [syros2025saga] proposed SAGA. SAGA makes users register agents with the central entity Provider and implement fine-grained interaction control using encrypted access control tokens, thereby balancing security and performance.
@@
D6: Defense for SEO Poisoning. To mitigate SEO poisoning, agent servers should deploy robust agent searching algorithms. For example, they can introduce adversarial training to enhance the model’s anti-manipulation ability, or conduct semantic blurring/replacing on search keywords, to prevent malicious agents from improving rankings. Besides, the search algorithms can deploy a random factor to ensure a ratio of randomly selected agents in the final list. Meanwhile, dynamically updating parameters and inducing historical response quality are also helpful.
(iii) Against Unique Risks in P2P-based Communication.
@@
D7: Task Lifecycle Monitoring. We think the non-convergence problem is stubborn and hard to eliminate as long as the P2P architecture is not changed fundamentally. As a result, the method of mitigating this problem is to monitor the task lifecycle. Each access point should deploy a coordinator. For agent-agent communication, this coordinator monitors the execution status. When it detects that the task interaction is trapped in a loop (e.g., no progress after N consecutive rounds of responses) or the communication time exceeds a threshold, it forcibly terminates the non-convergent communication. At the same time, the abnormal patterns and the communication participants are recorded for further analysis. He et al. [he2025attention] propose a Trust Management System (TMS), which deploys message-level and agent-level trust evaluation. TMS can dynamically monitor agent communication, execute threshold-driven filtering strategies, and achieve agent-level violation record tracking. Zhang et al. [zhang2025g] propose G-Memory, a hierarchical memory system. G-Memory manages the interaction history of agent communication through three-layer graph structures of Insight Graph, Query Graph, and Interaction Graph, thereby achieving the evolution of the agent team. Ebrahimi et al. [ebrahimi2025adversary] propose an anti-adversarial multi-agent system based on Credibility Score. It models query answering as an iterative cooperative game, distributes rewards through Contribution Score, and dynamically updates the credibility of each agent based on historical performance.
VI-C3 Defenses on Semantic Interpretation Layer (L3)
The defenses are classified based on the risks from L3.
(i) Against Universal Risks in all Communication Modes.
@@ VI-C3 Defenses on Semantic Interpretation Layer (L3)
D1: Cross-agent Input Detection. To mitigate cross-agent contamination, systems should enforce strict input isolation to prevent malicious or manipulative content from propagating across agents. Instead of directly concatenating raw messages, agents should extract structured, task-relevant information while filtering out control-oriented, emotional, or manipulative content that may induce cognitive deviation. Furthermore, deploying a safety coordination agent to review, sanitize, or flag inter-agent messages can effectively mitigate the potential attack propagation within multi-agent systems.
D2: Permission Classification and Control. To mitigate privacy leakage, the access control among agents is a core component for the future agent ecosystem. Although end-to-end encryption can avoid sniffing from external attackers to some extent, it cannot mitigate the unintentional privacy leakage among agents. Access control should assign access permission tags to different agents and ensure that agents need to attach permission proofs when communicating. In this way, agents with low-level permissions cannot obtain the high-level sensitive information from other agents. Zhang et al. [zhang2025llm] design the AgentSandbox framework, which uses the separation of persistent agents and temporary agents, data minimization, and I/O firewalls, realizing the security of agents in solving complex tasks. Kim et al. [kim2025prompt] propose the PFI framework, which defends against authority-related attacks through three major technologies: agent isolation, secure untrusted data processing, and privilege escalation guards. Wang et al. [wang2025agentspec] propose AgentSpec. It allows users to define rules containing trigger events, predicate checks, and execution mechanisms through a domain-specific language to ensure the safety of agent behavior.
D3: Malicious Instruction Filtering. To defend against description poisoning, multi-agent systems should authenticate and validate capability descriptions before they are consumed by the planner or other agents. Instead of relying on free-form, self-reported descriptions, systems should adopt structured and schema-constrained capability profiles, combined with cryptographic attestation or signatures to ensure their integrity. Additionally, cross-agent consistency checks and behavior-based verification can detect mismatches between an agent’s declared role and its actual behavior. A centralized or distributed semantic auditor can further sanitize or flag suspicious descriptions, preventing poisoned role definitions from misleading routing, delegation, or decision-making processes.
(ii) Against Unique Risks in CS-based Communication.
D4: Multi-source Semantic Cross-Validation. To mitigate centralized semantic poisoning, the semantic state maintained by the server must be protected through multi-layered verification and controlled update mechanisms. The system should enforce multi-source semantic cross-checking, where agent roles, capability descriptors, and trust scores are validated by multiple independent models or rule-based analyzers rather than a single server-side LLM, reducing the impact of prompt injection or metadata forgery. All semantic-state modifications should be recorded in an append-only audit log to prevent silent tampering and to provide traceability for administrative review. Finally, the server-side LLM responsible for semantic arbitration should be sandboxed to isolate transient prompt interactions from persistent semantic state, preventing injected content from contaminating long-term system semantics.
D5: Defense for Semantic Rewriting Attack. To defend against semantic rewriting attacks, CS-based orchestrators must ensure that all normalization and reformulation actions preserve the original task intent and safety semantics. This requires applying semantic differencing techniques that compare rewritten outputs with their originals using independent LLMs or symbolic analyzers to detect alterations in intent, constraints, or safety-critical content. The rewriting process must remain transparent and verifiable. First, rewriting rules should be explicit and auditable, aiming to remove ambiguity rather than introduce new semantics. Second, the system should expose both the original and rewritten messages to downstream agents and record them in an audit log to ensure traceability. Third, the server must follow strict policies during rewriting and avoid altering high-risk fields, such as safety requirements, tool-invocation parameters, or execution constraints. Finally, the system can attach semantic-integrity tags to rewritten messages, enabling downstream agents to verify that the rewrite preserves the intended meaning and to reject any messages whose semantic integrity cannot be confirmed.
(iii) Against Unique Risks in P2P-based Communication.
D6: Defense for Cognitive and Ethical Drift. To limit semantic and normative divergence in decentralized P2P systems, agents should perform cross-agent consistency checks before acting on received messages. Each agent can maintain a minimal shared semantic contract—such as standardized intent labels or schema-constrained reasoning outputs—that allows heterogeneous agents to align their interpretations. For ethically sensitive decisions, agents should execute rule-based constraint checks or lightweight justification verification to ensure their actions remain within globally agreed boundaries. Additionally, adopting small-scale consensus or multi-agent cross-validation for high-impact actions can prevent a misaligned agent from unilaterally steering group decisions. These mechanisms provide practical guardrails that reduce both cognitive drift and downstream ethical collapse in P2P collaboration.
D7: Defense for Contextual Fragmentation. To counter contextual fragmentation in P2P communication, agents must maintain a resilient approximation of shared conversational state despite operating with only local information. This can be achieved by exchanging compact state summaries, intent tokens, or hashed dialogue checkpoints that allow agents to partially reconstruct global context without requiring centralized control. Incorporating temporal validity checks and staleness detection prevents agents from acting on outdated assumptions. For multi-hop interactions, systems may employ context-reconciliation modules that merge or align divergent local states when agents reconnect. These mechanisms collectively preserve semantic coherence in asynchronous P2P environments and reduce misinterpretation caused by fragmented conversational history.
@@ VI-D Takeaways
In this section, we categorize two major agent-agent communication architectures: CS-based and P2P-based. Regardless of CS-based or P2P-based architecture, L1 needs to address data tampering and DoS attacks through end-to-end encryption and redundant transmission. Besides, L1 of CS-based architecture also requires additional protection of the centralized controller; L2 risks and defenses vary by architecture: CS-based needs to resist registration pollution and SEO poisoning through registration verification and load balancing, while P2P-based relies on distributed identity authentication to mitigate agent spoofing; L3 of CS-based needs to prevent centralized semantic poisoning and semantic rewriting attacks, while P2P-based addresses cognitive drift and contextual fragmentation through semantic contract alignment and context integration. The layered characteristic of the three-layered architecture makes risk prevention and control of different architectures more targeted, providing structured security guarantees for agent-agent communication.